The scores of 10 randomly selected students in a 50-point Statistics

Examination are as follows 49, 35, 38, 25, 34, 21, 47, 40, 35, 10


1
Expert's answer
2021-11-25T12:14:32-0500

"Average = \frac{Total}{Count}\\

Average = \frac{49+ 35+38+ 25+ 34+ 21+47+ 40+ 35+10}{10}\\
=\frac{334}{10}\\
=\:33.4"